Circular No. 6830

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998T IN IC 694 W.-d. Li, University of California at Berkeley (UCB), forwards the re-

port from C. Li that Zhou Wan has found an apparent supernova (unfiltered CCD mag 15.4) on Mar. 3 UT as part of the Beijing Astronomical Observatory Supernova Survey. As with SN 1998S (cf. IAUC 6829), this object was confirmed by the Lick Supernova Survey at unfiltered mag ∼ 14.5 on Mar. 4.3, with the corresponding KAIT position given as α = 11^{h}28^{m}32^{s}, δ = +58^{o}33′.7 (equinox 2000.0); SN 1998T appears in a dust lane of the irregular host galaxy that has no apparent nucleus, so the relative offsets cannot be accurately determined, and its close proximity to a knot makes it difficult to measure (the magnitudes were derived from subtracted images). Observations of the same field prior to Mar. 3 do not show an object at the position of SN 1998T.

A. V. Filippenko and E. C. Moran, UCB, report that on Mar. 5 they obtained a Keck-2 spectrogram (range 510–900 nm, resolution 1 nm) of the stellar-looking knot on the southern side of the dust lane of IC 694 (the eastern part of the irregular galaxy Arp 299 = Markarian 171, of which the western part is NGC 3690; see Wynn-Williams et al. 1991, Ap.J. 377, 426). SN 1998T is indeed a supernova, of type Ib; strong P-Cyg profiles of He I lines are visible. Narrow nebular emission lines are very intense.

SUPERNOVA 1998S IN NGC 3877 Filippenko and Moran also report: “A low-resolution Keck-2 spectrum,

obtained as above, shows that the type-II SN 1998S (cf. IAUC 6829) is peculiar; the spectrum contains prominent H and He emission lines superposed on a featureless blue continuum. In many respects, the spectrum resembles that of the type-IIn SN 1997eg (cf. IAUC 6794): the lines have multiple components, including a very narrow (unresolved) one that is probably produced by flash-ionized circumstellar gas. This object might become very luminous at radio and x-ray wavelengths, due to strong interaction between the ejecta and circumstellar material.”

V854 CENTAURI This R CrB-type variable is fading, as indicated by the following vi-

sual magnitude estimates, in part from the AAVSO: Feb. 4.67 UT, 7.6 (F. Farrell, Christies Beach, Australia); 21.9, 8.4 (B. Monard, Pretoria, South Africa); 25.22, 8.3 (A. Pereira, Cabo da Roca, Portugal); 26.96, 8.7 (Monard); Mar. 1.20, 9.1 (Pereira); 2.92, 9.0 (Monard); 3.85, 8.9 (Monard).

1998 March 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green